OpenStudy (destinyyyy):

Solve S = 4v2 for v. -4(4-v)= -2(2v-1) v-16+4v = -2(2v-1) v-16+4v = -4v + 2 -16+5v = -4v + 2 5v = -4v + 18 9v = 18 v = 2 Did i do this right??

OpenStudy (anonymous):

69

OpenStudy (destinyyyy):

What?

OpenStudy (acacia):

So Divide by 4 to get s/4=v^2 Take the square root to get: v=+/- sqrt(s)/2

OpenStudy (destinyyyy):

What?? O.o Sooo confused

OpenStudy (acacia):

S = 4v^2 We can do the same operation to both side of the equation and it will still be equal. We divide by 4.

OpenStudy (acacia):

S/4 = v^2 Now we take the root of both sides. sqr(S)/2 = v And that is the answer, if my guess was correct about v^2.

OpenStudy (destinyyyy):

Honestly i got this answer from someone else. But wasnt sure if they were correct because i dont know how to solve this problem at all. So i really do not understand what your saying. I just started doing algerbra 2 online and it didnt show how to solve a problem like this

OpenStudy (acacia):

Oh okay well the answer you got is correct

OpenStudy (destinyyyy):

okk.. One out of 8 done :/ I want to know how to solve these my self. But i cant find anything to show examples

OpenStudy (acacia):

What online program is it?

OpenStudy (destinyyyy):

Flvs

OpenStudy (acacia):

Oh I haven't heard of that one

OpenStudy (destinyyyy):

Its the main one that's used. Most schools recommend using it. I just started the class and i'm only on the second assignment and already confused